OpenStudy (anonymous):

Solve the following...... -8 = -10x-4y -4x= 16-8y

OpenStudy (rina.r):

USE SIMULATION OR ELIMATION METHOD TO SOVEL WHICH ONE WOULD U CHOOSE?

OpenStudy (rina.r):

WHICH ONE DO U WANT ME TO USE?

OpenStudy (anonymous):

@Rina.r Idk I just dont know how to solve it.

OpenStudy (anonymous):

@Rina.r Elemination

OpenStudy (rina.r):

ok let me show u.

OpenStudy (anonymous):

@Rina.r Ok

OpenStudy (rina.r):

i am doning on my note book so plz wait

OpenStudy (anonymous):

@Rina.r Ok

OpenStudy (rina.r):

-8=10x-4(-4x+16) -8=-10x-8x+10 -8-16=-10x-8x -24=-18x x=1.33 value -4x=16-8y -4(1.33)=16-8y -5.32=16-8y -8y=16+5.32 -8y=21.32 y=2.665.

OpenStudy (rina.r):

i am not sure about the answer but that's how it should be dont with susbituation mathod.